# Emergent spinon-holon Feshbach resonance in a doped Majumdar-Ghosh model

### v1

@@ II ARPES spectra
In Fig. S2, we show the interaction-dependent majority ($\sigma=\;\uparrow$, a-d) and minority ($\sigma=\;\downarrow$, e-h) ARPES spectra at fixed momentum $k=1.0$ at different magnetizations. We label branches according to their qualitative behavior and illustrate the fitting of the polaron branches. Here we also briefly discuss the spectral weights of the polaron and the molecular branches.
The minority molecule is not visible in the minority spectrum because of its vanishing spectral weight: We remove a $\downarrow$-spinon from a GS where practically all $\downarrow$-spinons are bound in singlets, see Fig. S2e-h. Conversely, the majority molecule ARPES weight increases for increasing $m$ since we have more free $\uparrow$-spinons in the GS. This is especially relevant for the majority spectrum where the entire majority molecule spectral weight originates from free $\uparrow$-spinons in the GS, see Fig. S2a-d
@@
Figure S2: ARPES spectra. We plot the majority ($\sigma=\;\uparrow$, a-d) and minority ($\sigma=\;\downarrow$, e-h) ARPES spectrum Eq. (3) of Hamiltonian (1) at fixed GS momentum $k=1.0$ for $L=81$ and at different magnetizations $m$. We label the branches according to their qualitative behavior. Fits of the polaron branches are marked in blue. We find an almost flat band for the attractive minority molecule in the inset of (d).

### v2

@@ II ARPES spectra
Figure 6: ARPES spectra. We plot the majority ($\sigma=\;\uparrow$, a-d) and minority ($\sigma=\;\downarrow$, e-h) ARPES spectrum Eq. (3) of Hamiltonian (1) at fixed GS momentum $k=1.0$ for $L=81$ and at different magnetizations $m$. We label the branches according to their qualitative behavior. We fit the polaron branches using the ansatz Eq. (5) and plot the fits in blue. We find an almost flat band for the attractive minority molecule in the inset of (d).
In Fig. 6, we show the interaction-dependent majority ($\sigma=\;\uparrow$, a-d) and minority ($\sigma=\;\downarrow$, e-h) ARPES spectra at fixed momentum $k=1.0$ at different magnetizations. We label branches according to their qualitative behavior and illustrate the fitting of the polaron branches (see below). Here we also briefly discuss the spectral weights of the polaron and the molecular branches.
The minority molecule is not visible in the minority spectrum because of its vanishing spectral weight: We remove a $\downarrow$-spinon from a GS where practically all $\downarrow$-spinons are bound in singlets, see Fig. 6e-h. Conversely, the majority molecule ARPES weight increases for increasing $m$ since we have more free $\uparrow$-spinons in the GS. This is especially relevant for the majority spectrum where the entire majority molecule spectral weight originates from free $\uparrow$-spinons in the GS, see Fig. 6a-d
@@
We set $J=1$ for convenience and fit the attractive/repulsive polaron branches using the ansatz
$E_{\mathrm{a/r}}(g,m)=\frac{A_{\mathrm{a/r}}}{g-g_{\mathrm{c,a/r}}}\times m+B_{\mathrm{a/r}}(m),$ (5)
where crucially $A_{\mathrm{a/r}}$ and $g_{\mathrm{c,a/r}}$ are independent of the magnetization. We find good agreement with $g_{\mathrm{c,a}}=0.2$, $g_{\mathrm{c,r}}=-1.3$, and $A_{\mathrm{a}}=A_{\mathrm{r}}=0.56$, the fits are the blue curves in Fig. 6a-c.
We can estimate the location of the Feshbach resonance as
$g_{\mathrm{c}}=\frac{g_{\mathrm{c,a}}+g_{\mathrm{c,r}}}{2}=-0.55,$ (6)
which is in good agreement with Fig. 2b. $g_{\mathrm{eff}}(g_{c})$ can be estimated from Eq. (2) as
$g_{\mathrm{eff}}(g_{c})$ =^$=\frac{\langle\hat{\mathcal{H}}_{\mathrm{int}}\rangle}{m\times\langle\hat{n}^{\mathrm{h}}\rangle}$
$=\frac{E_{\mathrm{r}}(g_{\mathrm{c}},m)-E_{\mathrm{r}}(g\to\infty,m)}{m\times\langle\hat{n}^{\mathrm{h}}\rangle}$
$\quad\;\,-\frac{E_{\mathrm{a}}(g_{\mathrm{c}},m)-E_{\mathrm{a}}(g\to-\infty,m)}{m\times\langle\hat{n}^{\mathrm{h}}\rangle}$
$=\frac{4A_{\mathrm{a}}}{\bigl{[}g_{\mathrm{c,a}}-g_{\mathrm{c,r}}\bigr{]}\times\langle\hat{n}^{\mathrm{h}}\rangle}$
$\approx 1.5\approx 2.7\times|g_{\mathrm{c}}|,$ (7)
thus confirming that the effective interaction we find is Feshbach-like, i.e. resonantly enhanced! Intuitively, the interaction energy corresponds to the bending of the polaron branches which is a direct consequence of the resonantly enhanced interaction.